Tohban Report 2011-07-13

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 06 July 2011
End Date: 13 July 2011
Tohban: Hazel Bain
Tohban email:
Next Tohban: Lindsay Glesener
List all reports


Solar Activity

The Sun was quiet this week with 7 C class flares. There are currently 6 AR's on disk. STEREO B shows one AR which will rotate onto the disk within the next few days.

How many GOES flares occurred?

 Flares above B, C, M, X class were     27     7     0     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were     13     5     0     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were     10     1     0     0

There were RHESSI flares/GOES flares 52 / 34 over the time range 06-Jul-11 13-Jul-11

Memory Management

We remained in normal/active decimation mode. Eclipse data was taken for both the front and rear detectors. There were a reduced number of passes due to THEMIS and shuttle passes. SSR is no longer emptying by the end of each pass set (min ~25%, max ~45%). Eclipse data will be reduced to 4 minutes at the start and end of night.

Data Gaps

There were no data gaps this week.

Spacecraft Status

Detector 6 voltage was lowered from 2990.2 to 2867.7 V at 19:37 UT July 6, 2011. No further arcs were detected for detector 6 this week. Detector 8 showed some noisy data (2011-07-06 13:55, 2011-07-08 04:30, 2011-07-10 04:45), Jim to check if these are bad data packets.

Personal tools